OpenStudy (anonymous):

Find the average rate of change for f(x) from -2 to 8 and the equation of the secant line containing (-2, f(-2)) and (8, f(8)). Please show all of your work.

OpenStudy (anonymous):

Find the average rate of change for f(x) = x^3 -6x^2 -5 from -2 to 8 and the equation of the secant line containing (-2, f(-2)) and (8, f(8)). Please show all of your work.

OpenStudy (anonymous):

\[\frac{f(8)-f(-2)}{8-(-2)}\] is the slope, aka \[\frac{f(8)-f(-2)}{10}\]

OpenStudy (saifoo.khan):

That's what i was writing^

OpenStudy (anonymous):

lol now it is an actual function, but the work is the same, just compute the numerator

OpenStudy (anonymous):

look at the question i just posted.. the first one was missing some numbers